OpenStudy (anonymous):

Which one of the following equations illustrates the commutative property? 8 + 3 = 4 + 7 If 3 + 5 = a, then a = 3 + 5 4 + x = x + 4 3(2 + 5) = 6 + 15

OpenStudy (anonymous):

Commutative Property Says that: \[\huge \color{green}{a + b = b+a}\] Now, according to this which one will be your answer??

OpenStudy (anonymous):

@waterineyes it would be 4+x=x+4 correct?

OpenStudy (anonymous):

Yes you are absolutely correct in figuring out it....

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Welcome dear..